# A perturbative non-Markovian treatment to low-temperature spin decoherence

### v1

@@ 1 Computational Details
We obtain hyperfine tensors for V1 - V4 from electronic structure calculations using the ORCA 6 quantum chemistry package. 29 Following previous work, we optimize the geometries using density functional theory with B3LYP and density fitting, with a def2-triple-$\zeta$ polarized basis for hydrogen, sulfur, carbon, and oxygen and a core-polarized CP(PPP) basis for vanadium. 7, 8, 20, 37, 38, 23, 5 The vanadium-oxygen bond is oriented to lie along the $z$-axis, and we compute hyperfine tensors from the optimized geometries and use the $zz$ component in Eq. 12 for each hydrogen spin. We include contributions to the hyperfine tensors from the dipolar coupling between the electron and nuclear spins along with the isotropic Fermi contact term. We include a bath of random hydrogen spins for comparison to experimental results. We constrain the average distance from the electron to the added hydrogen spins to be greater than 11 Å, and compute the additional interactions from the point-dipole approximation. The randomized coordinates of the extra hydrogen spins were generated once and used for each of V1 - V4.
@@ 2.2 Supporting Information
Details on the derivation of the TCL equation, the spin Hamiltonian, the error analysis, heteronuclear spin-pairs, and the molecular systems of interest.

### v2

@@ 1 Computational Details
We obtain hyperfine tensors for V1 - V4 from electronic structure calculations using the ORCA 6 quantum chemistry package. 35 Following previous work, we optimize the geometries using density functional theory with B3LYP and density fitting, with a def2-triple-$\zeta$ polarized basis for hydrogen, sulfur, carbon, and oxygen and a core-polarized CP(PPP) basis for vanadium. 10, 11, 24, 43, 44, 27, 7 The vanadium-oxygen bond is oriented to lie along the $z$-axis, and we compute hyperfine tensors from the optimized geometries and use the $zz$ component in Eqs11 and 14 for each hydrogen spin. We include contributions to the hyperfine tensors from the dipolar coupling between the electron and nuclear spins along with the isotropic Fermi contact term. We include a bath of random hydrogen spins for comparison to experimental results. To generate a physically reasonable bath, we place the molecule at the center of a cube with a 40 Å edge length. We then estimate the concentration of solvent molecules in the cube using the density of a common solvent molecule, N,N-dimethylformamide-d7 (DMF) with an isotopic purity of 99%, or 1% undeuterated hydrogens. We increase the concentration of the undeuterated hydrogens to account for the presence of undeuterated counter-ion molecules, tetraphenylphosphonium (Ph4P+) present in the experimental system. The interactions between the solvent hydrogen spins and electron spin are computed using the point-dipole approximation and added to Eqs. 11 and 14. The randomized coordinates of the extra hydrogen spins were generated once and used for each of V1 - V4.
@@ 2.2 Supporting Information
Details on the derivation of the TCL2 and TCL4 equations, the spin Hamiltonian, the error analysis, the factorization of spin pairs, comparison to the APPA, and molecular systems.